OpenStudy (s):

How to find the critical number of f(x)=(x^2-3x+2)^1/2

OpenStudy (anonymous):

\[f'(x)=\frac{2x-3}{2\sqrt{x^2-3x+2}}\] is a start

OpenStudy (anonymous):

set the numerator equal to zero and solve to find the critical number

OpenStudy (s):

why there is 2 in the front?

OpenStudy (anonymous):

the derivative of \[\sqrt{f(x)}\] is \[\frac{f'(x)}{2\sqrt{f(x)}}\] by the chain rule

OpenStudy (s):

correct